OpenStudy (anonymous):

find the derivative:

OpenStudy (anonymous):

\[y=x ^{x ^{2}}\]

OpenStudy (anonymous):

so far i have \[x ^{2}(x)^{x ^{2}-1}(1)\]

OpenStudy (anonymous):

dy/dx=x^2(x^(x^2-1) =

OpenStudy (anonymous):

thats what i have written already

OpenStudy (anonymous):

no x is a variable

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you are right thats the derivative

OpenStudy (anonymous):

so is the answer: \[x ^{x ^{2}+1}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[y = x ^{x ^{2}}=\exp(x ^{2}.\ln (x))\]

OpenStudy (anonymous):

and now you can derive

OpenStudy (anonymous):

why are you changing it to a natural log?

OpenStudy (anonymous):

it seems like youre making it more complicated than it has to be.

OpenStudy (anonymous):

when you have \[y=x ^{n} \]

OpenStudy (anonymous):

the condition to derive like that \[n.x ^{n-1}\]

OpenStudy (anonymous):

look up to see what i have already written down

OpenStudy (anonymous):

is that n is constant.

OpenStudy (anonymous):

yes I've seen what you've written

OpenStudy (anonymous):

and it's false because the power isn't constant

OpenStudy (anonymous):

if u take na log then u should also take nalog y then it become more complicated

OpenStudy (anonymous):

so how do i start the problem? take ln of both sides?

OpenStudy (anonymous):

so when you have something like that \[y = f(x)^{g(x)}\]

OpenStudy (anonymous):

\[lny=x^2 lnx\]

OpenStudy (anonymous):

you must first write it \[y = \exp ( g(x).\ln (f(x))\]

OpenStudy (anonymous):

is what i wrote juts above you right?

OpenStudy (anonymous):

no you don't have to do this ((\[lny=x ^{2}\ln(x)\]

OpenStudy (anonymous):

you know how to derive this \[\exp(f(x))\]

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\[(\exp(u))'=u'\exp(u)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for you u = x²ln(x) it's OK??

OpenStudy (anonymous):

u'= 2xln(x) + x³ . 1/x

OpenStudy (anonymous):

excuse it's \[u'=2xln(x) + x ^{2}\times \frac{ 1 }{x }\]

OpenStudy (anonymous):

you follow me ??